# Modeling Language as a Sequence of Thoughts

### v1

@@ Abstract
Transformer language models can generate strikingly natural text by modeling language as a sequence of tokens. Yet, by relying primarily on surface-level co-occurrence statistics, they fail to form globally consistent latent representations of entities and events, lack of which contributes to brittleness in relational direction (e.g., reversal curse), contextualization errors, and data inefficiency. On the other hand, cognitive science shows that human comprehension involves converting the input linguistic stream into compact, event-like representations that persist in memory while verbatim form is short-lived. Motivated by this view, we introduce Thought Gestalt (TG) model, a recurrent Transformer that models language at two levels of abstraction—tokens and sentence-level "thought" states. TG generates the tokens of one sentence at a time while cross-attending to a memory of prior sentence representations. In TG, token and sentence representations are generated using the same set of model parameters and trained with a single objective, the next-token cross-entropy: by retaining the computation graph of sentence representations written to memory, gradients from future token losses flow backward through cross-attention to optimize the parameters generating earlier sentence vectors. In scaling experiments, TG consistently improves efficiency over matched GPT-2 runs, among other baselines, with scaling fits indicating GPT-2 requires ~5–8% more data and ~33–42% more parameters to match TG’s loss. TG also reduces errors on relational direction generalization on a father-son reversal curse probe.
@@ 1 Introduction
Prior work in cognitive science suggests that in humans, language functions primarily as a transmission system for expressing underlying thoughts, rather than as the constitutive medium of cognition itself [1, 2, 3]. On this view, comprehension involves decoding the serial stream of language to construct mental models that encode temporal sequences, causal relations, and entity properties in a situation [4, 5]. These models organize information into coherent gestalts– latent representations whose properties are not reducible to the sum of their parts [6, 7, 8]. From this perspective, linguistic input is a lossy, serial code for these underlying structures, not the format in which mental models are stored.
Large language models (LLMs) have achieved remarkable capabilities by modeling language as a sequence of tokens and optimizing next-token prediction [9, 10, 11, 12]. However, standard Transformers represent context primarily as token embeddings tied to positional indices; consequently, the model’s internal state is inherently coupled to linear word order [13, 14, 15]. Moreover, because training is driven by surface-level distributional statistics—estimating which token is likely to follow a local pattern of surrounding tokens—models can generate fluent text without constructing a coherent, globally consistent situation model, which leads to systematic failures on compositional, multi-step, and long-context tasks [16, 2, 17, 18].
Several phenomena highlight the fragility of this token-centric approach. Lepori et al. [19] describe contextualization errors, where lower Transformer layers fail to resolve ambiguities because the representations of relevant prior context are not yet fully contextualized. Similarly, the reversal curse shows that models trained on "A is B" often fail to infer "B is A," treating the two directions of a relational fact as distinct statistical patterns rather than a unified semantic relation [20, 21]. Furthermore, LLMs remain strikingly data-inefficient: state-of-the-art models are pretrained on trillions of tokens [9, 12], exceeding the linguistic input of a human child (30 million words) by five orders of magnitude [22]. Yet, humans can often infer the meaning of a new word from a single exposure by linking it to existing conceptual structures [23, 24]. These limitations motivate architectures that build latent representations at a higher level of abstraction than tokens that encode knowledge and causal relations into robust "thought-level" states.
@@
In this work, we introduce the Thought Gestalt (TG) model, a recurrent Transformer architecture designed to model language at two levels of abstraction: tokens and sentence-level thoughts (Figure 1). Our approach is motivated by cognitive evidence that human perceivers segment continuous streams of information into discrete events to organize memory, where the exact verbatim form is short-lived and confined to a narrow span, while earlier content is retained as stable, high-level representations of events and relations [25, 26, 27]. To emulate this, TG predicts the tokens of one sentence at a time while updating and attending to an abstracted memory of sentence representations. Each sentence representation is a d_model vector that captures the sentence’s holistic meaning, or "gestalt," rather than merely summing its tokens. TG uses sentence boundaries as a structural proxy for thought boundaries. While not a one-to-one map to thought boundaries, sentence boundaries serve as natural cognitive junctures where background information is integrated and concepts are updated [28, 29]. Learning sentence representations is thus a first step toward building generative systems that can learn situation models and latent thought representations.
TG is
inspired by the Sentence Gestalt model [30], a recurrent network that incrementally maps a sequence of words onto a single distributed event representation—the “gestalt”—from which role–filler propositions can be decoded. In TG, the holistic vector representation of each sentence is extracted from the contextualized hidden state at the end-of-sentence (<EOS>) position and written into a recurrent memory that conditions subsequent token predictions. Crucially, TG optimizes sentence representations by maintaining their computation graph across memory writes, allowing next-token prediction losses on later sentences to backpropagate through cross-attention into the parameters that produced earlier sentence gestalts. To keep backpropagation tractable, we control the computation depth of memory representations with a simple curriculum over the number of sentences processed continuously without resetting the memory.
@@
This design distinguishes TG from recent architectures that learn sentence embeddings via auxiliary objectives, such as next-sentence prediction, token reconstruction, sentence ordering, or contrastive alignment, on top of a language-modeling loss [31, 32, 33, 34, 35]. Empirical analyses suggest such auxiliary losses can be brittle and even degrade downstream generalization [36, 37]. In contrast, TG employs no separate encoder or auxiliary sentence-level loss: token and sentence representations are generated using the same set of model parameters and are supervised solely via their contribution to the next-token prediction objective. Furthermore, unlike standard Transformers and Gisting [38], where summary tokens are processed in parallel with lexical tokens via attention masks—forcing lower layers to attend to noisy, unrefined representations of prior context—TG enables even early layers to attend to a memory of fully contextualized gestalts, directly targeting the contextualization errors identified by Lepori et al. [19].
@@
Empirically, TG yields systematic gains in both learning efficiency and representational robustness. TG is consistently more data and parameter efficient than matched GPT-2 baselines pretrained on WikiText-103. In dataset-scaling experiments (–M training tokens; M non-embedding parameters), TG achieves – lower test perplexity at every scale (e.g., vs. at M), corresponding to an effective – reduction in the training tokens needed to reach the same loss (Figure 2a). In parameter scaling experiments at fixed M tokens, TG preserves an advantage across a M–M model size range, where GPT-2 would require a – parameter increase to match TG (Figure 2b). Comparisons to alternative baselines—including adding sentence-boundary bias to GPT-2 via special tokens, replacing TG’s sentence-based recurrence with fixed token spans, and compression via attention masks (“gisting”)—show that TG’s gains hinge on conditioning token generation on semantically coherent, fully contextualized units which are diminished when replaced with arbitrary token chunks or context abstraction alone (Figure 5). Additionally, in a controlled father–son relation probe, TG improves relational direction generalization (measured by reversed-direction likelihood) substantially faster than GPT-2 (Figure 4). Finally, ablation experiments indicate that higher-capacity TG variants, with added self and cross attention within each layer, can further boost performance, and confirm that end-to-end training of sentence representations via retaining gradient flow through memory, is essential for the observed gains (Table 1).
2 Other Works
Learning sentence representations.
Large Concept Model (LCM) [44] similarly operates at the sentence level, but trains an autoregressive model to predict the next sentence embedding in a frozen SONAR representation space, which is then decoded into tokens, rather than directly predicting a token sequence. In contrast, the Thought Gestalt (TG) model does not use a frozen encoder or any separate sentence-level loss: token and sentence spaces are learned jointly using the same next-token prediction objective. This avoids brittle auxiliary losses while still yielding contextualized sentence embeddings that can be reused as high-level state of prior content.
Sequence compression and gisting.
In TG, sentence representations play the role of compression tokens: the model extracts a single gist vector per sentence from a mid–upper layer after full contextualization with within-sentence tokens and the sentence memory, and writes it into a recurrent memory without detaching computation graph. The only supervision for these gists comes from the same next-token losses on later sentences, backpropagated through the memory. Unlike Compressive Transformers, TG does not introduce a separate compression loss; and unlike standard gisting, the same fully contextualized sentence representation is reused in the cross-attention path at every layer (each Transformer layer applies its own learned projections to these vectors). Thus, even early layers can query a stable, contextualized summary of prior content rather than raw, noisy activations, helping alleviate long-range contextualization errors [19].
Recurrence and memory in Transformers.
Another design axis concerns extending Transformers to reuse past representations to enable accessing content beyond a fixed context window. Transformer-XL caches hidden states from previous segments and lets the current segment attend to them, providing segment-level recurrence but truncating gradients at the cached states [47]. Block-Recurrent Transformers apply a Transformer layer recurrently over blocks of tokens, with a recurrent state that each block attends to, combining local self-attention within a block with an RNN-like state that carries information across blocks [48]. Recurrent Memory Transformer adds dedicated memory tokens that are passed between segments and updated by self-attention, providing a differentiable memory channel across long sequences [49]. Memorizing Transformers, on the other hand, add an external key–value store of past activations that can be queried with kNN retrieval at inference time, extending the effective context while keeping the memory non-differentiable [50].
TG shares with these models the design principles of reusing past computation, but its recurrent state resides in an external differentiable memory of sentence-level gists rather than token-level caches or memory tokens encoding arbitrary token sequences, and these gists are trained only via next-token prediction loss at future sentence positions. This design ties recurrence, compression, and sentence structure together: long-range information is carried forward in compact sentence representations that are optimized end-to-end for generative performance. These semantically grounded units are natural candidates for storage and retrieval in a long-term memory system, in the spirit of the Memorizing Transformer but with sentence- or thought-aligned chunks that retrieval work suggests are more effective than arbitrary fixed-size token blocks [51, 52].
@@ 3 Main Design Principles
The Thought Gestalt (TG) model learns language as a sequence of sentence-level thoughts using a single transformer stack that functions both as a token decoder and a sentence encoder (Figure 3). Concretely, TG processes a document as a sequence of sentence steps. At step the model (i) predicts the tokens of sentence by interleaving causal self-attention over that sentence and cross-attention to a memory of earlier contextualized sentence vectors; then (ii) compresses into a single vector which is written to memory at the next iteration without detaching gradients. This section describes the TG architecture, its read/write interface to memory, and the training and data preparation pipelines that enable stable, scalable learning.
@@ 3.1 Data Preparation
Our data pipeline converts raw corpora into sentence-level training samples aligned with TG’s computation model. Specifically, we: (i) split datasets into document-level text units based on boundaries provided by the source (i.e., a single Wikipedia article); (ii) split documents into sentences; (iii) enclose each sentence with special boundary tokens; (iv) slice documents into continuous sentence streams to create training examples with bounded gradient-graph depth; and (v) construct batches by sampling uniformly from sentence streams, constrained by a target number of lexical tokens per batch, to stabilize GPU memory usage and optimization (detailed in Appendix C.1).
@@
We preprocess the training corpora by first separating each document (a standalone Wikipedia article) based on title formatting, and then splitting text into sentences using the “SaT Capped” method. SaT Capped predicts sentence boundaries at the token level and is robust to imperfect punctuation; it enforces a maximum sentence length by applying punctuation-aware fallback rules to split sentences longer than a given maximum token length (set to tokens in our experiments) into shorter, semantically coherent spans [53, 54, 55]. We selected this method as it achieves the highest AutoBLEU scores for sentence-level reconstruction in a similar architecture [55].
Sentence tokenization.
Following sentence splitting, each sentence is tokenized, padded to the set maximum length , and enclosed by special boundary markers: a start-of-sentence token (<BOS>) and an end-of-sentence token (<EOS>). The <BOS> token enables the TG model to predict the first lexical token of a new sentence; it sets the context to a non-empty state to enable cross-attending to the sentence memory. The <EOS> token marks the end of sentence generation and serves as the designated position for extracting the hidden state of sentence representation, (see §3.2). For the final sentence of a document, we append an additional <EOD> marker before the <EOS> to signal the end of text generation. The effective length of each input sentence tensor is thus fixed at (accounting for <BOS>, <EOD>, and <EOS>).
@@
TG retains the computation graph of sentence representations written to memory to learn sentence encodings via the next-token prediction loss. Although at any sentence step , the model only attends directly to the most recent sentence representations (where is the memory capacity), the computation graph for those stored representations includes their own cross-attention to earlier sentences. Consequently, the gradient dependency chain extends backwards recursively (e.g., ) to the first sentence processed after resetting the memory. Therefore, to keep training tractable and avoid storing computation graphs of unbounded depth, we slice training documents (not validation or test documents) into contiguous sentence streams of at most sentences. Each stream functions as an independent training example, and the memory is reset between streams. This strategy preserves long-range conditioning while capping the maximum depth of backpropagation. To maintain optimization stability given the variable number of lexical tokens per stream, batches are constructed by uniformly sampling streams until a target range of lexical tokens per batch is reached (details in Appendix C.1).
@@ 3.2 Model
Sentence representation and memory.
Let be the tokens of the current sentence ( is the fixed length of each processed sentence tensor) and let denote the layer- hidden states in the shared stack (model dimension ). Let index the end-of-sentence (EOS) token in the context. At a fixed mid layer (we use ), we form the sentence vector . Here denotes the sentence head, which is a single linear layer (depth 1). Choosing a mid layer for sentence representation extraction is motivated by prior evidence showing that intermediate transformer layers carry the most contextual and transferable features, whereas top layers increasingly specialize in token-level decisions (i.e., next-token lexicalization) [56, 18, 57].
TG maintains a rolling sentence memory of capacity . In our experiments, we set to approximate the 1024-token context window of the GPT-2 baseline (the average sentence length in our training corpus is 25 tokens). At step , the memory contains the most recent sentence vectors, ordered from oldest to most recent. At the end of step , the computed is appended to this memory, and if full, the oldest entry is removed.
Cross-attention to sentence memory.
Let and let the memory contain the most recent sentence vectors . We form cached key/value matrices
where denotes sinusoidal positional encodings (sentence-index positions) added to the keys only. At a cross-attention block , the queries are the layer- token hidden states projected as . The block forms projected keys/values and and computes
with per-head dimension . are computed once per sentence step and reused by all cross-attention layers when processing , while each block applies its own .
Injecting position information into keys—rather than values—decouples sequence order from semantic content. Prior work shows that this design improves robustness to sequence length variations by encoding position directly into attention scores [58, 59, 60]. In TG, token queries already carry token‑level positions from the input, and we inject sentence‑index positions only into the memory keys.
@@
We use a standard token embedding matrix and learned token positional embeddings ; the input to the model stack is . Each self-attention block is pre-norm, causal multi-head attention with residual connections. Within a block , token self-attention uses its own projections and standard scaled dot-product attention. (Regularization details are in Appendix C.2.)
@@
Each cross-attention block has a scalar, learnable memory gate that scales the cross-attention increment before it is added back via the residual path. These gates let the model modulate reliance on the memory pathway over training (e.g., up-weighting memory once sentence representations stabilize). See Appendix A for more detail.
@@
In standard Transformers, the prediction of the first token relies on a static <BOS> embedding, lacking local context. Because TG resets the self-attention window at every sentence boundary, this issue is exacerbated. To address this, we replace the static <BOS> embedding for sentence with the preceding sentence representation : This transforms the start token into a contextualized state. Consequently, the first lexical token of a new sentence can access prior context via two complementary pathways: (1) cross-attention to the memory, and (2) local self-attention to the -th position, which now explicitly encodes the previous sentence representation.

### v2

@@ Abstract
Transformer language models can generate strikingly natural text by modeling language as a sequence of tokens. Yet, by relying primarily on surface-level co-occurrence statistics, they fail to form globally consistent latent representations of entities and events, lack of which contributes to poor relational generalization (reversal curse), contextualization errors, and data inefficiency. On the other hand, cognitive science shows that human comprehension involves converting the input linguistic stream into compact, event-like representations that persist in memory while verbatim form is short-lived. Motivated by these cognitive findings, we introduce the Thought Gestalt (TG) Model, a recurrent transformer that models language at two levels of abstraction—tokens and sentence-level "thought" states. TG generates the tokens of one sentence at a time while cross-attending to a working memory of prior sentence representations. In TG, token and sentence representations are generated using a shared stack of transformer blocks and trained with a single objective, the next-token prediction loss: by retaining the computation graph of sentence representations written to the working memory, gradients from future token losses flow backward through cross-attention to optimize the parameters generating earlier sentence vectors. In scaling experiments, TG consistently improves data and parameter efficiency compared to matched GPT-2 runs, among other baselines, with scaling fits indicating GPT-2 requires ~5–8% more data and ~33–42% more parameters to match TG’s test loss. TG also reduces errors in relational-direction generalization on a father–son reversal curse probe.
@@ 1 Introduction
Prior work in cognitive science suggests that, in humans, language functions as a serial code for communicating underlying thoughts, rather than the constitutive medium of thought itself [54, 39, 16, 47]. On this view, comprehension involves decoding a linguistic stream to construct a situation model—a mental representation that encodes the temporal sequences, causal relations, and entities of the described event [59, 20]. Situation models support memory and later inference and are characterized as high-level conceptual representations integrated with prior knowledge rather than representations of the surface form of text [59, 39].
In contrast, modern Large Language Models (LLMs) learn by modeling language as a sequence of tokens and optimizing next-token prediction [46, 12, 1]. Although this has yielded models with remarkable fluency, a token-centric training signal can encourage brittle heuristics that capture surface-level statistical patterns rather than underlying concepts, leading to failures in generalization and compositional tasks [4, 15, 42]. The reversal curse is a concrete example: models trained on one relational direction (e.g., “A is B”) often fail to generalize to the inverse (“B is A”), treating the two directions of a relational fact as distinct patterns rather than a unified semantic encoding [5, 34]. While Berglund et al. [5] note that the reversal curse is not observed for in-context learning in their evaluated models (e.g., GPT-4), in §3.4 we show that a substantial in-context directional asymmetry exists in smaller models at the scale of GPT-2 when a relation is queried in the opposite direction of the prompt.
Another shortcoming of standard transformer models is the contextualization errors identified by Lepori et al. [32], where in lower transformer layers, later tokens attend to earlier token representations before they have been fully contextualized to resolve ambiguities. Finally, many modern models are trained on trillions of tokens [12], exceeding estimates of children’s language exposure by orders of magnitude (on the order of tens of millions of words [22]). These gaps motivate architectures that learn and reuse latent representations at a higher level of abstraction than tokens, i.e., models that organize information into coherent gestalts—holistic representations whose properties are not reducible to the sum of their parts [57]—that capture the underlying concepts and relations expressed by text [39, 20].
@@
In this work, we introduce the Thought Gestalt (TG) model, a recurrent transformer architecture designed to model language at two levels of abstraction: tokens and sentence-level thoughts. TG processes one sentence at a time (a "sentence step"), maintaining token-level information of only the current sentence, while context is maintained as a working memory of prior sentence representations: holistic representations, or "gestalts", each compressing an entire sentence into a vector (Figure 1). This design is motivated by cognitive evidence that humans segment continuous streams of text into discrete events to organize memory: the verbal form is short-lived and confined to a narrow span, while earlier content is retained as stable, high-level representations of events and relations [50, 24, 60]. TG is also inspired by the Sentence Gestalt model [55], which incrementally maps a word sequence onto a single event representation from which role–filler structure can be decoded. TG similarly uses sentence boundaries as a structural proxy for thought boundaries. While not a one-to-one mapping, sentence boundaries serve as natural cognitive junctures where background information is integrated and concepts are updated [21, 38]. Learning sentence representations is thus a first step toward building generative systems that can learn situation models and latent thought representations.
Architecturally, TG interleaves self-attention over tokens of the current sentence with cross-attention to a fixed-capacity working memory of prior sentence representations, where tokens attend causally within the current sentence and can reach prior sentences only via sentence-gestalt vectors (similar to gisting [44]). Each sentence representation is built from the contextualized hidden state at the end-of-sentence (<EOS>) token position (Figure 3). After each sentence step, the new sentence representation is appended to memory and the oldest entry is evicted when the memory is full. Crucially, TG retains the computation graph of sentence representations when writing to the working memory, allowing next-token prediction losses from later sentences to backpropagate through cross-attention into the parameters that produced earlier sentence gestalts (see Figure 1). Because sentence gestalts are contextualized summaries of semantically coherent chunks of prior content, they are natural candidates for caching and retrieval at inference time to enable continual learning (as the Memorizing transformer [58]); here we focus on their role as a differentiable context for pretraining.
@@
Our design of sentence representations learned end-to-end under the standard next-token objective distinguishes TG from recent architectures that learn sentence embeddings via auxiliary objectives, such as next-sentence prediction, token reconstruction, sentence ordering, or contrastive alignment, on top of a language-modeling loss [13, 14, 31, 18, 27]. Empirical analyses suggest such auxiliary losses can be brittle and even degrade downstream generalization [36, 2]. In contrast, TG employs no separate encoder or auxiliary sentence-level loss: token and sentence representations are generated using the same set of model parameters and are supervised solely via their contribution to the next-token prediction objective. Furthermore, unlike standard transformers, TG enables early layers to attend to a memory of fully contextualized sentence gestalts, directly targeting the contextualization errors identified by Lepori et al. [32].
@@
Empirically, TG yields systematic gains in both learning efficiency and representational robustness. TG is consistently more data and parameter efficient than matched GPT-2 baselines pretrained on WikiText-103 (Figure 2). In dataset-scaling experiments (–M training tokens; M non-embedding parameters), TG achieves – lower test perplexity at every scale (e.g., vs. at M), corresponding to an effective – reduction in the training tokens needed to reach the same loss. In parameter scaling experiments at fixed M tokens, TG preserves an advantage across a M–M model size range, where GPT-2 would require – more parameters to match TG. Comparisons to alternative baselines-including adding sentence-boundary bias to GPT-2 via special tokens, replacing TG’s sentence-based recurrence with fixed token spans, and sentence-level gisting-style attention masking-show that TG’s gains depend on conditioning token generation on semantically coherent, fully contextualized units; performance degrades when sentences are replaced with arbitrary token blocks or when access to prior tokens is restricted through compression tokens without recurrence (Figure 4). Additionally, in an in-context father–son probe, TG improves relational-direction generalization substantially faster than GPT-2, as measured by the likelihood of the target token when the completion query reverses the direction of the prompt (Figure 5). Finally, ablations confirm that the working memory of sentence gestalts and retaining gradient flow through memory are essential for TG’s performance and that adding per-layer self and cross attention capacity further improves performance (Table 1).
@@ 2 Main Design Principles
The Thought Gestalt (TG) model learns language as a sequence of sentence-level thoughts using a single transformer stack that functions both as a token decoder and a sentence encoder (Figure 3). Concretely, TG processes a document as a sequence of sentence steps. At sentence step , where is the number of sentences in the sentence stream, TG (i) predicts the tokens of sentence by interleaving causal self-attention over its tokens and cross-attention to a working memory of earlier contextualized sentence vectors; and (ii) compresses sentence into a single vector, , which is written to memory without detaching gradients. This section describes the TG architecture, its read/write interface to the working memory, and the training and data preparation pipelines that enable stable, scalable learning. For details on batch construction and stochastic regularization, refer to Appendix B; for ablations of key TG components see §3.5, and for training-time analyses of memory use and gradient flow see Appendix C and Appendix D.
@@ 2.1 Data Preparation
Our data pipeline converts raw corpora into sentence-level training samples aligned with TG’s computation model. Specifically, we: (i) split datasets into document-level text units based on boundaries provided by the source (i.e., a single Wikipedia article); (ii) split documents into sentences; (iii) enclose each sentence with special boundary tokens; (iv) slice documents into continuous sentence streams of fixed maximum length to create training examples with bounded gradient-graph depth; and (v) construct batches by sampling uniformly from sentence streams, constrained by a target number of lexical tokens per batch, to stabilize GPU memory usage and optimization (detailed in Appendix B.1).
@@
We preprocess the training corpora by first separating each document (a standalone Wikipedia article) based on title formatting, and then splitting text into sentences using the “SaT Capped” method. SaT Capped predicts sentence boundaries at the token level and is robust to imperfect punctuation; it enforces a maximum sentence length by applying punctuation-aware fallback rules to split sentences longer than a given maximum token length (set to tokens in our experiments) into shorter, semantically coherent spans [17, 41, 3]. We selected this method as it achieves the highest AutoBLEU scores for sentence-level reconstruction, as reported in the Large Concept Model study [3].
Sentence tensorization and boundary markers.
After sentence splitting, each sentence is tokenized (tokenizer/vocabulary shared across all baselines) and padded to the set maximum sentence length tokens. We then form sentence tensors by adding explicit boundary markers: a start-of-sentence token <BOS> and an end-of-sentence token <EOS>. The <BOS> token enables the TG model to predict the first lexical token of a new sentence; it sets the context to a non-empty state to enable cross-attending to the working memory. The <EOS> token marks the end of sentence generation and serves as the designated position for extracting the hidden state used to form the sentence representation, (see §2.2). In addition, we reserve a dedicated slot immediately before <EOS> for an end-of-document marker <EOD>. This slot is only added for the final sentence of a document to signal the end of text generation and is <PAD> for all non-final sentences. Thus, every sentence tensor has the same fixed total length
@@
TG retains the computation graph of sentence representations written to memory to learn sentence encodings via the next-token prediction loss. At sentence step , the model cross-attends only to the most recent sentence vectors in the forward pass. However, because the computation graph of memory entries is retained, the computation graph of stored sentence vectors retain links to the graph of earlier sentences they cross-attended to when they were formed. As a result, gradients from step can recursively backpropagate into the computation graph of older sentence representations that are no longer present in the working memory (Appendix Fig. 6). Therefore, to keep training tractable and avoid storing computation graphs of unbounded depth, we slice training documents (not validation or test documents) into contiguous sentence streams with a maximum length cap of sentences. Each stream functions as an independent training example, and the memory is reset between streams, which truncates the gradient chain. This strategy preserves long-range conditioning while capping the maximum depth of backpropagation. To maintain optimization stability given the variable number of lexical tokens per stream, batches are constructed by uniformly sampling streams until a target range of lexical tokens per batch is reached (details in Appendix B.1).
@@ 2.2 Model
Sentence representation.
Let be the tokens of the current sentence, denote the hidden states at layer of the model stack (with ), and index the <EOS> position. At a fixed mid layer ( in our setting), we form the sentence vector
Here denotes the sentence head, a single linear layer (depth 1). At the end of step , the computed is appended to the working memory, without detaching its compute graph, and if working memory is full, the oldest entry is removed. Importantly, lives in the same -dimensional space as token hidden states ( in our experiments). Choosing a mid layer to extract the hidden state used for generating the sentence representation, , is motivated by prior evidence showing that intermediate transformer layers carry the most contextual and transferable features, whereas top layers increasingly specialize in token-level decisions (i.e., next-token lexicalization) [19, 35, 53].
Cross-attention to working memory.
TG maintains a rolling working memory of fixed capacity . In our experiments, we set to approximate the 1024-token context window of the GPT-2 baseline (the average sentence length in our training corpus is 25 tokens). At step , the working memory contains the most recent sentence representations , where , ordered from oldest to most recent. For efficiency, we form (once per sentence step) the unprojected memory key/value matrices, which are reused by all cross-attention layers when processing sentence
where are sinusoidal sentence-index positional encodings added to keys only. Each cross-attention block has its own learned projections, thus at cross-attention layer we use
Multi-head cross-attention is then computed per head (with and ) and concatenated in the standard way; in compact form,
Injecting sentence order information into keys (rather than values) decouples positional information from semantic content and encodes position directly into attention scores, which prior work shows can improve robustness under length variation [48, 49, 43]. In TG, token queries already include token-level positional information via , thus we add positional structure only to the memory keys. Appendix D provides evidence that, as pretraining scale increases, a larger fraction of the training signal flows through the cross-attention pathway, consistent with the analysis Appendix D showing that the model learns to rely increasingly on the sentence-memory as training progresses and sentence representations stabilize.
@@
We use a standard token embedding matrix and learned token positional embeddings ; the input to the model stack is . Each self-attention layer is pre-norm, causal multi-head attention with residual connections. Within a block , token self-attention uses its own projections and standard scaled dot-product attention.
@@
Each cross-attention layer has a scalar, learnable memory gate that scales the cross-attention increment before it is added back via the residual path. These gates let the model modulate reliance on the memory pathway over training (up-weighting memory once sentence representations stabilize (Appendix C)).
@@
In standard transformers, the prediction of the first token relies on a static <BOS> embedding, lacking local context. Because TG resets the context window at every sentence boundary, this issue is exacerbated. To address this, we replace the static <BOS> embedding for step with the preceding sentence representation , if available:
This transforms the start token into a contextualized state. Consequently, the first lexical token of a new sentence can access prior context via two complementary pathways: (1) cross-attention to the memory, and (2) local self-attention to the <BOS> position, which now explicitly encodes the previous sentence representation.